OpenStudy (anonymous):

Hi, i have a real analysis question: Using the definition , show that the following sets are closed E = {xeR:absolute value(5x-3)= absolute value(2x+4)} E={xeR: 1

OpenStudy (anonymous):

\[\left| 5x-3 \right|=\left| 2x+4 \right|\] and \[1\le x^{2}\le3\]